# Mixed FEM for poroelasticity with stress-dependent permeabilityKhan, Lamichhane, Ruiz-Baier & Villa-Fuentes

### v1

@@ 4.2 Analysis of the discrete problem
(4.11) $\widetilde{C}^{*}\big{(}1+\,\gamma\,C_{1}^{*}\big{)}\big{(}\|g\|_{0,\Omega}+\|p_{\mathrm{D}}\|_{1/2,\Gamma_{\mathrm{D}}}+\|\boldsymbol{u}_{\mathrm{D}}\|_{1/2,\Gamma_{\mathrm{D}}}+\|\boldsymbol{f}\|_{0,\Omega}\big{)}+\dfrac{\widetilde{C}^{*}}{c_{a}}\,\gamma^{2}r\leq r,$
where $\widetilde{C}^{*}$ and $C_{1}^{*}$ (depending on $c_{a}$, $\mu$, $\kappa_{1}$, $\kappa_{2}$, $C_{\widetilde{a}}$, $\beta^{*}$, $\widetilde{\beta}^{*}$) are the discrete versions of the constants $\widetilde{C}$ and $C_{1}$ (cf. (3.7) and (3.9)). Finally, it is clear that $(\boldsymbol{\sigma}_{h},\boldsymbol{u}_{h},\boldsymbol{\rho}_{h},{\boldsymbol{\varphi}}_{h},p_{h})$ is a solution to (4.6) if and only if $p_{h}$ satisfies
(4.12) $\mathbf{T}_{h}(p_{h})=p_{h}.$
@@
(4.13) $\frac{2\mathop{\rm max}\limits\{1,\,\kappa_{2}\}}{\min\{\widetilde{\gamma},\,\kappa_{1}\,,r\}}\,\bigg{\{}\widetilde{C}^{*}(1+C_{1}^{*}\gamma)(\|g\|_{0,\Omega}+\|p_{\mathrm{D}}\|_{1/2,\Gamma_{\mathrm{D}}}+\|\boldsymbol{u}_{\mathrm{D}}\|_{1/2,\Gamma_{\mathrm{D}}}+\|\boldsymbol{f}\|_{0,\Omega})+\dfrac{\gamma^{2}}{c_{a}}\biggl{(}\frac{1}{\kappa_{2}}+\widetilde{C}^{*}\,r\biggr{)}\bigg{\}}<1.$
Then, $\mathbf{T}_{h}$ (cf. (4.10)) has a unique fixed point $p_{h}\in\mathrm{W}_{h}$. Equivalently, problem (4.6) has a unique solution $(\boldsymbol{\sigma}_{h},\boldsymbol{u}_{h},\boldsymbol{\rho}_{h},{\boldsymbol{\varphi}}_{h},p_{h})\in\mathbb{H}^{\boldsymbol{\sigma}}_{h}\times\mathbf{H}^{\boldsymbol{u}}_{h}\times\mathbb{H}^{\boldsymbol{\rho}}_{h}\times\mathbf{H}^{{\boldsymbol{\varphi}}}_{h}\times\mathrm{W}_{h}$.In addition, there exists $C^{*}>0$, such that
(4.14) $\|\boldsymbol{\sigma}_{h}\|_{\mathbf{div};\Omega}+\|\boldsymbol{u}_{h}\|_{0,\Omega}+\|\boldsymbol{\rho}_{h}\|_{0,\Omega}+\|{\boldsymbol{\varphi}}_{h}\|_{\mathop{\rm div}\nolimits;\Omega}+\|p_{h}\|_{0,\Omega}\leq C^{*}(\|g\|_{0,\Omega}+\|p_{\mathrm{D}}\|_{1/2,\Gamma_{\mathrm{D}}}+\|\boldsymbol{u}_{\mathrm{D}}\|_{1/2,\Gamma_{\mathrm{D}}}+\|\boldsymbol{f}\|_{0,\Omega}+\gamma\,r\,).$
Proof 4.2.
First, we observe that, similar to the continuous case (as seen in the proof of Theorem 3.12), assumption (4.13) ensures the well-definedness of $\mathbf{T}_{h}$ and that $\mathbf{T}_{h}(\mathrm{W}_{h})\subseteq\mathrm{W}_{h}$. Now, by adapting the arguments used in Section 3.3 (cf. Lemmas 3.8 and 3.10), one can derive the following estimates
@@
$\|S_{2,h}(\widehat{\boldsymbol{\sigma}}_{1},\,\widehat{p}_{1})-S_{2,h}(\widehat{\boldsymbol{\sigma}}_{2},\,\widehat{p}_{2})\|_{0,\Omega}$ $\leq\frac{2\kappa_{2}\,\widetilde{C}^{*}}{\min\{\widetilde{\gamma},\,\kappa_{1}\}}\,\big{(}\|g\|_{0,\Omega}+\|p_{\mathrm{D}}\|_{1/2,\Gamma_{\mathrm{D}}}+\gamma\|\widehat{\boldsymbol{\sigma}}_{2}\|_{\mathbf{div};\Omega}\big{)}\|\widehat{p}_{1}-\widehat{p}_{2}\|_{0,\Omega}$
$\quad+\frac{2}{\min\{\widetilde{\gamma},\,\kappa_{1}\}}\,\gamma\,\|\widehat{\boldsymbol{\sigma}}_{1}-\widehat{\boldsymbol{\sigma}}_{2}\|_{\mathbf{div};\Omega},$
for all $\widehat{p}_{1},\,\widehat{p}_{2}\in\mathrm{W}_{h}$ and $\widehat{\boldsymbol{\sigma}}_{1},\,\widehat{\boldsymbol{\sigma}}_{2}\in\mathbb{H}^{\boldsymbol{\sigma}}_{h}$, which together with the definition of $\mathbf{T}_{h}$ (see (4.10)), yield
$\|\mathbf{T}_{h}(\widehat{p}_{1})-\mathbf{T}_{h}(\widehat{p}_{2})\|_{0,\Omega}$ $\leq\dfrac{2}{\min\{\widetilde{\gamma},\,\kappa_{1}\}}\,\bigg{\{}\kappa_{2}\,\widetilde{C}^{*}(1+C_{1}^{*}\gamma)(\|g\|_{0,\Omega}+\|p_{\mathrm{D}}\|_{1/2,\Gamma_{\mathrm{D}}}+\|\boldsymbol{u}_{\mathrm{D}}\|_{1/2,\Gamma_{\mathrm{D}}}+\|\boldsymbol{f}\|_{0,\Omega})$
@@
for all $\widehat{p}_{1},\,\widehat{p}_{2}\in\mathrm{W}_{h}$. In this way, using estimate (4.13) we obtain that $\mathbf{T}_{h}$ is a contraction mapping on $\mathrm{W}_{h}$, thus problem (4.12), or equivalently (4.6) is well-posed. Finally, analogously to the proof of Theorem  3.12 (see also Lemmas 3.8 and 3.10) we can obtain (4.14), which concludes the proof.
@@ 5.1 Preliminaries
Let the tuples $(\boldsymbol{\sigma},\boldsymbol{u},\boldsymbol{\rho},{\boldsymbol{\varphi}},p)\in\mathbb{H}_{\mathrm{N}}(\mathbf{div};\Omega)\times\mathbf{L}^{2}(\Omega)\times\mathbb{L}_{\mathrm{skew}}^{2}(\Omega)\times\mathbf{H}_{\mathrm{N}}(\mathop{\rm div}\nolimits;\Omega)\times\mathrm{L}^{2}(\Omega)$ and $(\boldsymbol{\sigma}_{h},\boldsymbol{u}_{h},\boldsymbol{\rho}_{h},{\boldsymbol{\varphi}}_{h},p_{h})\in\mathbb{H}^{\boldsymbol{\sigma}}_{h}\times\mathbf{H}^{\boldsymbol{u}}_{h}\times\mathbb{H}^{\boldsymbol{\rho}}_{h}\times\mathbf{H}^{{\boldsymbol{\varphi}}}_{h}\times\mathrm{H}^{p}_{h}$ be the unique solutions of (3.3) and (4.9), respectively.
Let us write $\texttt{e}_{\boldsymbol{\sigma}}=\boldsymbol{\sigma}-\boldsymbol{\sigma}_{h}$, $\texttt{e}_{\boldsymbol{u}}=\boldsymbol{u}-\boldsymbol{u}_{h}$, $\texttt{e}_{\boldsymbol{\rho}}=\boldsymbol{\rho}-\boldsymbol{\rho}_{h}$, $\texttt{e}_{{\boldsymbol{\varphi}}}={\boldsymbol{\varphi}}-{\boldsymbol{\varphi}}_{h}$ and $\texttt{e}_{p}=p-p_{h}$. As usual, for a given $(\widehat{\boldsymbol{\tau}}_{h},(\widehat{\boldsymbol{v}}_{h},\widehat{{\boldsymbol{\eta}}}_{h}))\in\mathbb{H}^{\boldsymbol{\sigma}}_{h}\times(\mathbf{H}^{\boldsymbol{u}}_{h}\times\mathbb{H}^{\boldsymbol{\rho}}_{h})$ and $(\widehat{{\boldsymbol{\psi}}}_{h},\widehat{q}_{h})\in\mathbf{H}^{{\boldsymbol{\varphi}}}_{h}\times\mathrm{H}^{p}_{h}$, we shall then decompose these errors into
(5.1) $\texttt{e}_{\boldsymbol{\sigma}}=\mbox{\boldmath{$\xi$}}_{\boldsymbol{\sigma}}+{\boldsymbol{\chi}}_{\boldsymbol{\sigma}},\qquad\texttt{e}_{\boldsymbol{u}}=\mbox{\boldmath{$\xi$}}_{\boldsymbol{u}}+{\boldsymbol{\chi}}_{\boldsymbol{u}},\qquad\texttt{e}_{\boldsymbol{\rho}}=\mbox{\boldmath{$\xi$}}_{\boldsymbol{\rho}}+{\boldsymbol{\chi}}_{\boldsymbol{\rho}},\qquad\texttt{e}_{{\boldsymbol{\varphi}}}=\mbox{\boldmath{$\xi$}}_{{\boldsymbol{\varphi}}}+{\boldsymbol{\chi}}_{{\boldsymbol{\varphi}}},\qquad\texttt{e}_{p}=\xi_{p}+\chi_{p},$
[…]
@@
$\begin{array}[]{rllll}\displaystyle a(\boldsymbol{\sigma},\boldsymbol{\tau})&+\quad\displaystyle b(\boldsymbol{\tau},(\boldsymbol{u},\boldsymbol{\rho}))&=\displaystyle H(\boldsymbol{\tau})-c(\boldsymbol{\tau},p)&\forall\,\boldsymbol{\tau}\in\mathbb{H}_{\mathrm{N}}(\mathbf{div};\Omega),\\[4.30554pt] \displaystyle b(\boldsymbol{\sigma},(\boldsymbol{v},{\boldsymbol{\eta}}))&&=\displaystyle F(\boldsymbol{v},{\boldsymbol{\eta}})&\forall\,\boldsymbol{v}\in\mathbf{L}^{2}(\Omega),\,\forall\,{\boldsymbol{\eta}}\in\mathbb{L}_{\mathrm{skew}}^{2}(\Omega),\end{array}$
@@
$\begin{array}[]{rllll}\displaystyle a(\boldsymbol{\sigma}_{h},\boldsymbol{\tau}_{h})&+\quad\displaystyle b(\boldsymbol{\tau}_{h},(\boldsymbol{u}_{h},\boldsymbol{\rho}_{h}))&=\displaystyle H(\boldsymbol{\tau}_{h})-c(\boldsymbol{\tau}_{h},p_{h})&\forall\,\boldsymbol{\tau}_{h}\in\mathbb{H}^{\boldsymbol{\sigma}}_{h},\\[4.30554pt] \displaystyle b(\boldsymbol{\sigma}_{h},(\boldsymbol{v}_{h},{\boldsymbol{\eta}}_{h}))&&=\displaystyle F(\boldsymbol{v}_{h},{\boldsymbol{\eta}}_{h})&\forall(\boldsymbol{v}_{h},{\boldsymbol{\eta}}_{h})\in\mathbf{H}^{\boldsymbol{u}}_{h}\times\mathbb{H}^{\boldsymbol{\rho}}_{h}.\end{array}$
From these relations we can obtain that for all $(\boldsymbol{\tau}_{h},(\boldsymbol{v}_{h},\boldsymbol{\rho}_{h}))\in\mathbb{H}^{\boldsymbol{\sigma}}_{h}\times(\mathbf{H}^{\boldsymbol{u}}_{h}\times\mathbb{H}^{\boldsymbol{\rho}}_{h})$, there holds
$\begin{array}[]{rllll}\displaystyle a(\texttt{e}_{\boldsymbol{\sigma}},\boldsymbol{\tau}_{h})&+\quad\displaystyle b(\boldsymbol{\tau}_{h},(\texttt{e}_{\boldsymbol{u}},\texttt{e}_{\boldsymbol{\rho}}))&=\displaystyle-c(\boldsymbol{\tau}_{h},\texttt{e}_{p}),\\[4.30554pt] \displaystyle b(\texttt{e}_{\boldsymbol{\sigma}},(\boldsymbol{v}_{h},{\boldsymbol{\eta}}_{h}))&&=\displaystyle 0\end{array}$
@@
(5.2) $\begin{array}[]{ll}\displaystyle a({\boldsymbol{\chi}}_{\boldsymbol{\sigma}},\boldsymbol{\tau}_{h})+b(\boldsymbol{\tau}_{h},({\boldsymbol{\chi}}_{\boldsymbol{u}},{\boldsymbol{\chi}}_{\boldsymbol{\rho}}))+b({\boldsymbol{\chi}}_{\boldsymbol{\sigma}},(\boldsymbol{v}_{h},{\boldsymbol{\eta}}_{h}))\\ \displaystyle\qquad=-a(\mbox{\boldmath{$\xi$}}_{\boldsymbol{\sigma}},\boldsymbol{\tau}_{h})-b(\boldsymbol{\tau}_{h},(\mbox{\boldmath{$\xi$}}_{\boldsymbol{u}},\mbox{\boldmath{$\xi$}}_{\boldsymbol{\rho}}))-b(\mbox{\boldmath{$\xi$}}_{\boldsymbol{\sigma}},(\boldsymbol{v}_{h},{\boldsymbol{\eta}}_{h}))-c(\boldsymbol{\tau}_{h},{\boldsymbol{\chi}}_{p})-c(\boldsymbol{\tau}_{h},\mbox{\boldmath{$\xi$}}_{p})\end{array}$
for all $(\boldsymbol{\tau}_{h},(\boldsymbol{v}_{h},\boldsymbol{\rho}_{h}))\in\mathbb{H}^{\boldsymbol{\sigma}}_{h}\times(\mathbf{H}^{\boldsymbol{u}}_{h}\times\mathbb{H}^{\boldsymbol{\rho}}_{h})$.
@@
$\begin{array}[]{rllll}\displaystyle\widetilde{a}_{\boldsymbol{\sigma},p}({\boldsymbol{\varphi}},{\boldsymbol{\psi}})&+\quad\displaystyle\widetilde{b}({\boldsymbol{\psi}},p)&=\displaystyle\widetilde{H}({\boldsymbol{\psi}})&\forall\,{\boldsymbol{\psi}}\in\mathbf{H}_{\mathrm{N}}(\mathop{\rm div}\nolimits;\Omega),\\[4.30554pt] \displaystyle\widetilde{b}({\boldsymbol{\varphi}},q)&-\quad\widetilde{c}(p,q)&=\displaystyle G(q)+c(\boldsymbol{\sigma},q)&\forall\,q\in\mathrm{L}^{2}(\Omega).\end{array}$
@@
$\begin{array}[]{rllll}\displaystyle\widetilde{a}_{\boldsymbol{\sigma}_{h},p_{h}}({\boldsymbol{\varphi}}_{h},{\boldsymbol{\psi}}_{h})&+\quad\displaystyle\widetilde{b}({\boldsymbol{\psi}}_{h},p_{h})&=\displaystyle\widetilde{H}({\boldsymbol{\psi}}_{h})&\forall\,{\boldsymbol{\psi}}_{h}\in\mathbf{H}^{{\boldsymbol{\varphi}}}_{h},\\[4.30554pt] \displaystyle\widetilde{b}({\boldsymbol{\varphi}}_{h},q_{h})&-\quad\widetilde{c}(p_{h},q_{h})&=\displaystyle G(q_{h})+c(\boldsymbol{\sigma}_{h},q_{h})&\forall\,q_{h}\in\mathrm{H}^{p}_{h}.\end{array}$
Then, using arguments similar to those in Lemma 3.10, by adding ±$\pm\widetilde{a}_{\boldsymbol{\sigma}_{h},p_{h}}({\boldsymbol{\varphi}},{\boldsymbol{\psi}}_{h})$, we have
$\begin{array}[]{rllll}\displaystyle\widetilde{a}_{\boldsymbol{\sigma}_{h},p_{h}}(\texttt{e}_{{\boldsymbol{\varphi}}_{h}},{\boldsymbol{\psi}}_{h})+\widetilde{b}({\boldsymbol{\psi}}_{h},\texttt{e}_{p_{h}})+\widetilde{b}(\texttt{e}_{{\boldsymbol{\varphi}}_{h}},q_{h})-\widetilde{c}(\texttt{e}_{p_{h}},q_{h})=-\int_{\Omega}\Big{(}\kappa(\boldsymbol{\sigma},p)^{-1}-\kappa(\boldsymbol{\sigma}_{h},p_{h})^{-1}\Big{)}\,{\boldsymbol{\varphi}}\cdot{\boldsymbol{\psi}}_{h}+c(\texttt{e}_{\boldsymbol{\sigma}},q_{h}),\end{array}$

### v2

@@ 4.2 Analysis of the discrete problem
$\widetilde{C}^{*}\big{(}1+\,\gamma\,C_{1}^{*}\big{)}\big{(}\|g\|_{0,\Omega}+\|p_{\mathrm{D}}\|_{1/2,\Gamma_{\mathrm{D}}}+\|{\bm{u}}_{\mathrm{D}}\|_{1/2,\Gamma_{\mathrm{D}}}+\|\bm{f}\|_{0,\Omega}\big{)}+\dfrac{\widetilde{C}^{*}}{c_{a}}\,\gamma^{2}r\leq r,$ (4.9)
where $\widetilde{C}^{*}$ and $C_{1}^{*}$ (depending on $c_{a}$, $\mu$, $\kappa_{1}$, $\kappa_{2}$, $C_{\widetilde{a}}$, $\beta^{*}$, $\widetilde{\beta}^{*}$) are the discrete versions of the constants $\widetilde{C}$ and $C_{1}$ (cf. (3.7) and (3.9)). Finally, it is clear that $(\bm{\sigma}_{h},{\bm{u}}_{h},\bm{\rho}_{h},{\bm{\varphi}}_{h},p_{h})$ is a solution to (4.5) if and only if $p_{h}$ satisfies
$\mathbf{T}_{h}(p_{h})=p_{h}.$ (4.10)
@@
$\frac{2\mathop{\rm max}\limits\{1,\,\kappa_{2}\}}{\min\{\widetilde{\gamma},\,\kappa_{1}\,,r\}}\,\bigg{\{}\widetilde{C}^{*}(1+C_{1}^{*}\gamma)(\|g\|_{0,\Omega}+\|p_{\mathrm{D}}\|_{1/2,\Gamma_{\mathrm{D}}}+\|{\bm{u}}_{\mathrm{D}}\|_{1/2,\Gamma_{\mathrm{D}}}+\|\bm{f}\|_{0,\Omega})+\dfrac{\gamma^{2}}{c_{a}}\biggl{(}\frac{1}{\kappa_{2}}+\widetilde{C}^{*}\,r\biggr{)}\bigg{\}}<1.$ (4.11)
Then, $\mathbf{T}_{h}$ (cf. (4.8)) has a unique fixed point $p_{h}\in\mathrm{W}_{h}$. Equivalently, problem (4.5) has a unique solution $(\bm{\sigma}_{h},{\bm{u}}_{h},\bm{\rho}_{h},{\bm{\varphi}}_{h},p_{h})\in\mathbb{H}^{\bm{\sigma}}_{h}\times\mathbf{H}^{{\bm{u}}}_{h}\times\mathbb{H}^{\bm{\rho}}_{h}\times\mathbf{H}^{{\bm{\varphi}}}_{h}\times\mathrm{W}_{h}$.In addition, there exists $C^{*}>0$, such that
$\displaystyle\|\bm{\sigma}_{h}\|_{\mathbf{div};\Omega}+\|{\bm{u}}_{h}\|_{0,\Omega}+\|\bm{\rho}_{h}\|_{0,\Omega}+\|{\bm{\varphi}}_{h}\|_{\mathop{\rm div}\nolimits;\Omega}+\|p_{h}\|_{0,\Omega}$
$\qquad\leq C^{*}(\|g\|_{0,\Omega}+\|p_{\mathrm{D}}\|_{1/2,\Gamma_{\mathrm{D}}}+\|{\bm{u}}_{\mathrm{D}}\|_{1/2,\Gamma_{\mathrm{D}}}+\|\bm{f}\|_{0,\Omega}+\gamma\,r\,).$ (4.12)
Proof.
First, we observe that, similar to the continuous case (as seen in the proof of Theorem 3.8), assumption (4.11) ensures the well-definedness of $\mathbf{T}_{h}$ and that $\mathbf{T}_{h}(\mathrm{W}_{h})\subseteq\mathrm{W}_{h}$. Now, by adapting the arguments used in Section 3.3 (cf. Lemmas 3.6 and 3.7), one can derive the following estimates
@@
$\|S_{2,h}(\widehat{\bm{\sigma}}_{1},\,\widehat{p}_{1})-S_{2,h}(\widehat{\bm{\sigma}}_{2},\,\widehat{p}_{2})\|_{0,\Omega}$ $\leq\frac{2\kappa_{2}\,\widetilde{C}^{*}}{\min\{\widetilde{\gamma},\,\kappa_{1}\}}\,\big{(}\|g\|_{0,\Omega}+\|p_{\mathrm{D}}\|_{1/2,\Gamma_{\mathrm{D}}}+\gamma\|\widehat{\bm{\sigma}}_{2}\|_{\mathbf{div};\Omega}\big{)}\|\widehat{p}_{1}-\widehat{p}_{2}\|_{0,\Omega}$
$\quad+\frac{2}{\min\{\widetilde{\gamma},\,\kappa_{1}\}}\,\gamma\,\|\widehat{\bm{\sigma}}_{1}-\widehat{\bm{\sigma}}_{2}\|_{\mathbf{div};\Omega},$
for all $\widehat{p}_{1},\,\widehat{p}_{2}\in\mathrm{W}_{h}$ and $\widehat{\bm{\sigma}}_{1},\,\widehat{\bm{\sigma}}_{2}\in\mathbb{H}^{\bm{\sigma}}_{h}$, which together with the definition of $\mathbf{T}_{h}$ (see (4.8)), yield
$\|\mathbf{T}_{h}(\widehat{p}_{1})-\mathbf{T}_{h}(\widehat{p}_{2})\|_{0,\Omega}$ $\leq\dfrac{2}{\min\{\widetilde{\gamma},\,\kappa_{1}\}}\bigg{\{}\kappa_{2}\,\widetilde{C}^{*}(1+C_{1}^{*}\gamma)(\|g\|_{0,\Omega}+\|p_{\mathrm{D}}\|_{1/2,\Gamma_{\mathrm{D}}}+\|{\bm{u}}_{\mathrm{D}}\|_{1/2,\Gamma_{\mathrm{D}}}+\|\bm{f}\|_{0,\Omega})$
@@
for all $\widehat{p}_{1},\,\widehat{p}_{2}\in\mathrm{W}_{h}$. In this way, using estimate (4.11) we obtain that $\mathbf{T}_{h}$ is a contraction mapping on $\mathrm{W}_{h}$, thus problem (4.10), or equivalently (4.5) is well-posed. Finally, analogously to the proof of Theorem  3.8 (see also Lemmas 3.6 and 3.7) we can obtain (4.1), which concludes the proof. ∎
@@ 5.1 Preliminaries
Let the tuples $(\bm{\sigma},{\bm{u}},\bm{\rho},{\bm{\varphi}},p)\in\mathbb{H}_{\mathrm{N}}(\mathbf{div};\Omega)\times\mathbf{L}^{2}(\Omega)\times\mathbb{L}_{\mathrm{skew}}^{2}(\Omega)\times\mathbf{H}_{\mathrm{N}}(\mathop{\rm div}\nolimits;\Omega)\times\mathrm{L}^{2}(\Omega)$ and $(\bm{\sigma}_{h},{\bm{u}}_{h},\bm{\rho}_{h},{\bm{\varphi}}_{h},p_{h})\in\mathbb{H}^{\bm{\sigma}}_{h}\times\mathbf{H}^{{\bm{u}}}_{h}\times\mathbb{H}^{\bm{\rho}}_{h}\times\mathbf{H}^{{\bm{\varphi}}}_{h}\times\mathrm{H}^{p}_{h}$ be the unique solutions of (3.3) and (4.7), respectively.
Let us write $\texttt{e}_{\bm{\sigma}}=\bm{\sigma}-\bm{\sigma}_{h}$, $\texttt{e}_{{\bm{u}}}={\bm{u}}-{\bm{u}}_{h}$, $\texttt{e}_{\bm{\rho}}=\bm{\rho}-\bm{\rho}_{h}$, $\texttt{e}_{{\bm{\varphi}}}={\bm{\varphi}}-{\bm{\varphi}}_{h}$ and $\texttt{e}_{p}=p-p_{h}$. As usual, for a given $(\widehat{\bm{\tau}}_{h},(\widehat{\mathbf{v}}_{h},\widehat{{\bm{\eta}}}_{h}))\in\mathbb{H}^{\bm{\sigma}}_{h}\times(\mathbf{H}^{{\bm{u}}}_{h}\times\mathbb{H}^{\bm{\rho}}_{h})$ and $(\widehat{{\bm{\psi}}}_{h},\widehat{q}_{h})\in\mathbf{H}^{{\bm{\varphi}}}_{h}\times\mathrm{H}^{p}_{h}$, we shall then decompose these errors into
$\texttt{e}_{\bm{\sigma}}=\bm{\xi}_{\bm{\sigma}}+{\bm{\chi}}_{\bm{\sigma}},\qquad\texttt{e}_{{\bm{u}}}=\bm{\xi}_{{\bm{u}}}+{\bm{\chi}}_{{\bm{u}}},\qquad\texttt{e}_{\bm{\rho}}=\bm{\xi}_{\bm{\rho}}+{\bm{\chi}}_{\bm{\rho}},\qquad\texttt{e}_{{\bm{\varphi}}}=\bm{\xi}_{{\bm{\varphi}}}+{\bm{\chi}}_{{\bm{\varphi}}},\qquad\texttt{e}_{p}=\xi_{p}+\chi_{p},$ (5.1)
with $\bm{\xi}_{\bm{\sigma}}:=\bm{\sigma}-\widehat{\bm{\tau}}_{h}$, ${\bm{\chi}}_{\bm{\sigma}}:=\widehat{\bm{\tau}}_{h}-\bm{\sigma}_{h}$, $\bm{\xi}_{{\bm{u}}}:={\bm{u}}-\widehat{\mathbf{v}}_{h},$ ${\bm{\chi}}_{{\bm{u}}}:=\widehat{\mathbf{v}}_{h}-{\bm{u}}_{h}$, $\bm{\xi}_{\bm{\rho}}:=\bm{\rho}-\widehat{{\bm{\eta}}}_{h}$, ${\bm{\chi}}_{\bm{\rho}}:=\widehat{{\bm{\eta}}}_{h}-\bm{\rho}_{h}$, $\bm{\xi}_{{\bm{\varphi}}}:={\bm{\varphi}}-\widehat{{\bm{\psi}}}_{h}$, ${\bm{\chi}}_{{\bm{\varphi}}}:=\widehat{{\bm{\psi}}}_{h}-{\bm{\varphi}}_{h}$, $\xi_{p}:=p-\widehat{q}_{h}$, and $\chi_{p}:=\widehat{q}_{h}-p_{h}$.
@@
$\begin{array}[]{rllll}\displaystyle a(\bm{\sigma},\bm{\tau})&+\quad\displaystyle b(\bm{\tau},({\bm{u}},\bm{\rho}))&=\displaystyle H(\bm{\tau})-c(\bm{\tau},p)&\forall\,\bm{\tau}\in\mathbb{H}_{\mathrm{N}}(\mathbf{div};\Omega),\\[4.30554pt] \displaystyle b(\bm{\sigma},(\mathbf{v},{\bm{\eta}}))&&=\displaystyle F(\mathbf{v},{\bm{\eta}})&\forall\,\mathbf{v}\in\mathbf{L}^{2}(\Omega),\,\forall\,{\bm{\eta}}\in\mathbb{L}_{\mathrm{skew}}^{2}(\Omega),\end{array}$
@@
$\begin{array}[]{rllll}\displaystyle a(\bm{\sigma}_{h},\bm{\tau}_{h})&+\quad\displaystyle b(\bm{\tau}_{h},({\bm{u}}_{h},\bm{\rho}_{h}))&=\displaystyle H(\bm{\tau}_{h})-c(\bm{\tau}_{h},p_{h})&\forall\,\bm{\tau}_{h}\in\mathbb{H}^{\bm{\sigma}}_{h},\\[4.30554pt] \displaystyle b(\bm{\sigma}_{h},(\mathbf{v}_{h},{\bm{\eta}}_{h}))&&=\displaystyle F(\mathbf{v}_{h},{\bm{\eta}}_{h})&\forall(\mathbf{v}_{h},{\bm{\eta}}_{h})\in\mathbf{H}^{{\bm{u}}}_{h}\times\mathbb{H}^{\bm{\rho}}_{h}.\end{array}$
From these relations we can obtain that for all $(\bm{\tau}_{h},(\mathbf{v}_{h},\bm{\rho}_{h}))\in\mathbb{H}^{\bm{\sigma}}_{h}\times(\mathbf{H}^{{\bm{u}}}_{h}\times\mathbb{H}^{\bm{\rho}}_{h})$, there holds
$\begin{array}[]{rllll}\displaystyle a(\texttt{e}_{\bm{\sigma}},\bm{\tau}_{h})&+\quad\displaystyle b(\bm{\tau}_{h},(\texttt{e}_{{\bm{u}}},\texttt{e}_{\bm{\rho}}))&=\displaystyle-c(\bm{\tau}_{h},\texttt{e}_{p}),\\[4.30554pt] \displaystyle b(\texttt{e}_{\bm{\sigma}},(\mathbf{v}_{h},{\bm{\eta}}_{h}))&&=\displaystyle 0,\end{array}$
@@
$\begin{array}[]{ll}\displaystyle a({\bm{\chi}}_{\bm{\sigma}},\bm{\tau}_{h})+b(\bm{\tau}_{h},({\bm{\chi}}_{{\bm{u}}},{\bm{\chi}}_{\bm{\rho}}))+b({\bm{\chi}}_{\bm{\sigma}},(\mathbf{v}_{h},{\bm{\eta}}_{h}))\\ \displaystyle\qquad=-a(\bm{\xi}_{\bm{\sigma}},\bm{\tau}_{h})-b(\bm{\tau}_{h},(\bm{\xi}_{{\bm{u}}},\bm{\xi}_{\bm{\rho}}))-b(\bm{\xi}_{\bm{\sigma}},(\mathbf{v}_{h},{\bm{\eta}}_{h}))-c(\bm{\tau}_{h},{\bm{\chi}}_{p})-c(\bm{\tau}_{h},\bm{\xi}_{p}),\end{array}$ (5.2)
for all $(\bm{\tau}_{h},(\mathbf{v}_{h},\bm{\rho}_{h}))\in\mathbb{H}^{\bm{\sigma}}_{h}\times(\mathbf{H}^{{\bm{u}}}_{h}\times\mathbb{H}^{\bm{\rho}}_{h})$.
@@
$\begin{array}[]{rllll}\displaystyle\widetilde{a}_{\bm{\sigma},p}({\bm{\varphi}},{\bm{\psi}})&+\quad\displaystyle\widetilde{b}({\bm{\psi}},p)&=\displaystyle\widetilde{H}({\bm{\psi}})&\forall\,{\bm{\psi}}\in\mathbf{H}_{\mathrm{N}}(\mathop{\rm div}\nolimits;\Omega),\\[4.30554pt] \displaystyle\widetilde{b}({\bm{\varphi}},q)&-\quad\widetilde{c}(p,q)&=\displaystyle{\tilde{F}}(q)+c(\bm{\sigma},q)&\forall\,q\in\mathrm{L}^{2}(\Omega).\end{array}$
@@
$\begin{array}[]{rllll}\displaystyle\widetilde{a}_{\bm{\sigma}_{h},p_{h}}({\bm{\varphi}}_{h},{\bm{\psi}}_{h})&+\quad\displaystyle\widetilde{b}({\bm{\psi}}_{h},p_{h})&=\displaystyle\widetilde{H}({\bm{\psi}}_{h})&\forall\,{\bm{\psi}}_{h}\in\mathbf{H}^{{\bm{\varphi}}}_{h},\\[4.30554pt] \displaystyle\widetilde{b}({\bm{\varphi}}_{h},q_{h})&-\quad\widetilde{c}(p_{h},q_{h})&=\displaystyle{\tilde{F}}(q_{h})+c(\bm{\sigma}_{h},q_{h})&\forall\,q_{h}\in\mathrm{H}^{p}_{h}.\end{array}$
Then, using arguments similar to those in Lemma 3.7, by adding ±$\pm\widetilde{a}_{\bm{\sigma}_{h},p_{h}}({\bm{\varphi}},{\bm{\psi}}_{h})$, we have
$\displaystyle\widetilde{a}_{\bm{\sigma}_{h},p_{h}}(\texttt{e}_{{\bm{\varphi}}_{h}},{\bm{\psi}}_{h})+\widetilde{b}({\bm{\psi}}_{h},\texttt{e}_{p_{h}})+\widetilde{b}(\texttt{e}_{{\bm{\varphi}}_{h}},q_{h})-\widetilde{c}(\texttt{e}_{p_{h}},q_{h})$
$\qquad=-\int_{\Omega}\Big{(}{\bm{\kappa}}(\bm{\sigma},p)^{-1}-{\bm{\kappa}}(\bm{\sigma}_{h},p_{h})^{-1}\Big{)}\,{\bm{\varphi}}\cdot{\bm{\psi}}_{h}+c(\texttt{e}_{\bm{\sigma}},q_{h}),$